OpenStudy (anonymous):

in how many ways can 6 boys be arranged in a line so that 3 particular boys may never come together?

OpenStudy (anonymous):

they can be first third and fifth or second fourth and sixth, and there are \(3!\) choice for each , and then \(3!\) choices for the other three slots so \(2\times 3!\times 3!\)

OpenStudy (anonymous):

@kropot72 look reasonable? just off the top of my head

OpenStudy (kropot72):

@satellite73 Looks exactly correct! Very good reasoning :)

OpenStudy (anonymous):

But I did not get some part :( Why you multiplied it with 2?? @satellite73

OpenStudy (anonymous):

When there are no restrictions the 6 boys can be arranged in a row in 6! = 720 way When the 3 boys are treated as one "unit", there are 4 persons and these 4 persons can be arranged in 4 ! = 24 ways The unit boys can be arranged among themselves in 3 ! = 6 ways. The number of ways the 3 boys sit together is 4! × 3!= 144 Now, there are only 2 possibilities: the boys are together or they are not. So the number of ways of arranging them so that the 3 boys are NOT together is: 720 - 144 = 576

OpenStudy (anonymous):

So who is right??

OpenStudy (anonymous):

i once got a similar problem, and the answer was also biggere tham i thought.

OpenStudy (kropot72):

There are two ways of interpreting this question: 1. The three boys that may never come together cannot come together as a trio and cannot come together as a pair. 2. The three boys that may never come together cannot come together as a trio but can come together as a pair. The solution by @satellite73 follows interpretation #1 and the solution by @Luis_Rivera follows interpretation #2.